OpenStudy (anonymous):

a tennis ball is shot vertically upward from a launcher with an initial velocity of 64 ft per second. When will the ball return to the ground?

OpenStudy (anonymous):

where you given an initial equation h(t) = ??

OpenStudy (anonymous):

what you are looking for is h(0)

OpenStudy (anonymous):

ahhhmmm.. that was the given question but there are choices...

OpenStudy (anonymous):

a. 4 seconds b. 4.8 sec c. 48 sec d. 48.5 sec

OpenStudy (anonymous):

using the information in the previous problem , how far off the ground will the ball be after 1 sec

OpenStudy (anonymous):

what is 64/16...do not have cal next to me

OpenStudy (whpalmer4):

Come on, 64/16 is 4, you don't need a calculator for that!

OpenStudy (anonymous):

so a. 4 sec is the answer?

OpenStudy (anonymous):

No it's the wrong question...(sarcasm)

OpenStudy (anonymous):

but where did you get the 16???

OpenStudy (anonymous):

thanks whpalmer.....I think so keneisha...:) basically the formula is H(t) = -16T^2 + 64T solve for H(0)

OpenStudy (anonymous):

From Physics we know that the ball will reach a height of h feet after t seconds where h and t are related at the formula, based on the 64 given...

OpenStudy (anonymous):

t is 0 at launch and 4 when it hits the ground...

OpenStudy (anonymous):

using the information in the previous problem , how far off the ground will the ball be after 1 sec

OpenStudy (anonymous):

use the same formula and t = 1

OpenStudy (anonymous):

a. 4.8 ft b. 48 ft c. 48.5 ft d. 485 ft

OpenStudy (whpalmer4):

\[h(t) = v_0t - \frac{1}{2}gt^2\]\[v_0 = 64 ft/s\]\[g = 32 ft/s^2\] \(v_0t\) gives you the upward travel, \(-\frac{1}{2}gt^2\) gives you the downward component due to gravity. So you set h(t) = 0 and solve for the two values of t. The first value (t = 0) is the launch, the second is the landing.

OpenStudy (anonymous):

waht will be the answer in the second question?

OpenStudy (whpalmer4):

Evaluate h(1) with the formula.

OpenStudy (anonymous):

i'm not really good in math...

OpenStudy (whpalmer4):

\[h(1) = 64(1) - \frac{1}{2}(32)(1)^2 = \]

OpenStudy (whpalmer4):

No time like the present to work on changing that!

OpenStudy (anonymous):

48 ft

OpenStudy (anonymous):

thanks a lot for the both of you you two are really good in mathematics

OpenStudy (whpalmer4):

There you go! And if you look at my graph, you'll see that at the point over the 1 on the x-axis, it's just about to 50 on the y-axis.